# Contrastive Diffusion Guidance for Spatial Inverse Problems

### v1

@@ Abstract
We consider the inverse problem of reconstructing the spatial layout of a place, a home floorplan for example, from a user’s movements inside that layout. Direct inversion is ill-posed since many floorplans can explain the same movement trajectories. We adopt a diffusion-based posterior sampler to generate layouts consistent with the measurements. While active research is in progress on generative inverse solvers, we find that the forward operator in our problem poses new challenges. The path planning process inside a floorplan is a non-invertible, non-differentiable function, and causes instability while optimizing using the likelihood score. We break-away from existing approaches and reformulate the likelihood score in a smoother embedding space. The embedding space is trained with a contrastive loss which brings compatible floorplans and trajectories close to each other, while pushing mismatched pairs far apart. We show that a surrogate form of the likelihood score in this embedding space is a valid approximation of the true likelihood score, making it possible to steer the denoising process towards the posterior. Across extensive experiments, our model CoGuide produces more consistent floorplans from trajectories, and is more robust than differentiable-planner baselines and guided-diffusion methods.
@@ 1 Introduction
Consider a user walking around in her home for a few minutes. Using some sensor, e.g., a smartphone, the user’s trajectory has been recorded. This trajectory is a sequence of location measurements inside the home, $\mathbf{y}=[y_{1},y_{2},\dots y_{n}]$, along which the user has walked. We ask, given this trajectory measurement, is it possible to infer the floorplan $\mathbf{x}$ of the home, where floorplan is the dimensions and layouts of the walls in the home. Observe that this is a spatial inverse problem, modeled as $\mathbf{y}=\mathcal{A}(\mathbf{x},\mathbf{n})$, because the way the human user walks, $\mathbf{y}$, is indeed a function of the layout of the home $\mathbf{x}$. This function is the $\mathcal{A}(.)$ operator, a policy in the user’s brain that plans the path from point A to point B, for a given floorplan $\mathbf{x}$. This path-planner is complex—it models for factors such as distance walked, collision with walls and furniture, time to walk, number of turns, etc. Optimizations with such functions inherit this complexity since tiny changes in the floorplan—say a small hole in one wall—can drastically change the planned path. Hence, $\mathcal{A}(.)$ is in the regime of non-linear, non-differentiable, and partially observable operators, presenting a relatively new question (to the best of our knowledge) to diffusion-based inverse solvers. This paper concentrates on this specific problem of inverse floorplan estimation, but shows the potential to extend the approach to a broader family of spatial inverse problems.
@@
We train our Diffusion prior using public floorplan datasets; during inference, our method CoGuide generates floorplans for given (sparse, medium, or dense) trajectories. Results reliably outperform $6$ different baselines: $3$ that are augmentations of DPS with path-planners Yonetani et al. (2021); Kirilenko et al. (2023); Liu et al. (2024), $2$ that are established inverse solvers Zhu et al. (2023); Wang et al. (2024), and $1$ classifier free guidance (CFG) diffusion model Dhariwal & Nichol (2021) that (over)fits to the joint distribution of trajectories and floorplans. We believe CoGuide has potential beyond this specific application of floorplan inference; we discuss early thoughts on generalization and follow-on research directions to leverage contrastive learning in diffusion-based inverse solvers.
@@ 2 Preliminaries
Diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020); Song et al. (2020a; b) are a class of generative models capable of producing high-quality samples across a wide range of domains, including images Dhariwal & Nichol (2021), audio Kong et al. (2020); Liu et al. (2023), video Ho et al. (2022), and 3D data Luo & Hu (2021); Poole et al. (2022). These models define generation as the reversal of a forward noising process, formalized using a stochastic differential equation (SDE).
@@
Here, the prior score is easy to approximate by a trained diffusion model $s_{\theta}(\mathbf{x}_{t},t)$. The likelihood score, in contrast, is intractable as $\mathbf{y}$ depends only on $\mathbf{x}_{0}$, not directly on $\mathbf{x}_{t}$. In response to this, prior work such as Diffusion Posterior Sampling (DPS) Chung et al. (2022) have approximated the likelihood term as $p_{t}(\mathbf{y}|\mathbf{x}_{t})\approx p(\mathbf{y}|\hat{\mathbf{x}}_{0})$ where $\hat{\mathbf{x}}_{0}:=\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})=\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$ is obtained from a single denoising step using Tweedie’s formula: $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\mathbf{x}_{t}+\sigma_{t}^{2}\nabla_{\mathbf{x}_{t}}\log\;p_{t}(\mathbf{x}_{t})$ This formulation has been successfully used to guide diffusion models in solving a variety of inverse problems across scientific domains Zheng et al. (2025). CoGuide builds on this framework but makes a necessary departure owing to the challenges posed by the path-planning operator. We briefly discuss these operators before formulating the floorplan inference inverse problem.
Path Planners as Forward Operators. Unlike typical inverse problems (e.g., deblurring or inpainting), our measurements $\mathbf{y}$ are human-walked trajectories; thus, the forward operator $\mathcal{A}$ encodes a walking policy through an indoor layout (the floorplan $\mathbf{x}$). Directly modeling human navigation is difficult, so we use path-planners as a proxies LaValle (2006); Hart et al. (1968); Gammell et al. (2015); Noreen et al. (2016). Empirical evidence indicates people favor short, direct routes with few turns Tong & Bode (2022) which aligns well with shortest-path planning. Therefore, the classical A* algorithm Hart et al. (1968) is a good choice for the forward operator. The floorplan is discretized into a grid graph, and given a start and end location, A* computes the shortest collision-free path, which we take as the predicted walking trajectory. To fit into the inverse problem framework, differentiable variants of A*, such as Neural A* Yonetani et al. (2021), TransPath Kirilenko et al. (2023), NRRT Wang et al. (2020), and Takahashi et al. (2019) are of interest; they all aim to enable gradient-based learning. Recent diffusion-based planners (DiPPeR Liu et al. (2024), PbDiff Luo et al. (2024)) further provide differentiable path generation. When inserted as forward operators, these planners often induce non-smooth objectives where small layout perturbations can cause large path changes, which in turn hinders convergence in gradient-based optimization. We analyze causes for this instability next and motivate our embedding-space likelihood surrogate.
@@ 3.1 Problem Formulation
Fig. 1 (Left) top row shows an unknown 2D floorplan $\mathbf{x}\in\mathbb{R}^{m\times n}$. A user walks in this floorplan (from point A to a destination point B, then from B to another destination C, and so on) and records a sequence of location measurements. The union of all these location measurements on a 2D image give us the trajectories $\mathbf{y}\in\mathbb{R}^{m\times n}$ as shown in Fig. 1 (Left) bottom row. Given location sensors are noisy, the forward process is $\mathbf{y}=\mathcal{A}(\mathbf{x})+\mathbf{n}$ where $\mathcal{A}:\mathbb{R}^{m\times n}\!\to\!\mathbb{R}^{m\times n}$ approximates the human walk using a (generally nonlinear) A* path-planner, and $\mathbf{n}$ models additive Gaussian noise present in the location sensor.
@@
Approximating the likelihood requires propagating $\hat{\mathbf{x}}_{0}$ through the $\mathcal{A}(.)$ operator and using it’s gradient to steer the diffusion prior $s_{\theta}(\mathbf{x}_{t},t)$; doing this stably lies at the heart of our problem.
@@ 3.2 Guidance through Contrastive Learning
where $\mathbf{x}\in\mathcal{X}$ and $\mathbf{y}\in\mathcal{Y}$ such that $\|f_{\varphi}(\mathbf{x})\|_{2}=\|g_{\psi}(\mathbf{y})\|_{2}=1$. We then use these encoders to define the likelihood surrogate for a pair of inputs $(\mathbf{x},\mathbf{y})$ in the form of an un-normalized distribution:
$\pi(\mathbf{y}|\mathbf{x})\propto\exp(\langle f_{\varphi}(\mathbf{x}),\;g_{\psi}(\mathbf{y})\rangle\,/\,\tau)$ (7)
@@
Contrastive similarity as a likelihood surrogate. To see why this is a valid approach to approximating the true likelihood, we note that InfoNCE links contrastive learning to density estimation. Following Contrastive Predictive Coding Oord et al. (2018) we observe that the optimal contrastive classifier recovers a likelihood ratio, i.e., its logit approximates $\log\;p(\mathbf{y}|\mathbf{x})-\log\;p(\mathbf{y})$ up to an additive constant. That is, when the InfoNCE loss attains its optimum, we get
$\exp(\langle f_{\varphi}(\mathbf{x}),g_{\psi}(\mathbf{y})\rangle\,/\,\tau)$ $\propto\frac{p(\mathbf{y}|\mathbf{x})}{p(\mathbf{y})}\implies\frac{1}{\tau}\langle f_{\varphi}(\mathbf{x}),\;g_{\psi}(\mathbf{y})\rangle=\log\;p(\mathbf{y}|\mathbf{x})-\log\;p(\mathbf{y})+C$
where $C$ is a constant independent of $\mathbf{x}$. Taking gradients with respect to $\mathbf{x}$ exactly recovers the likelihood score on the right-hand side i.e., $\frac{1}{\tau}\nabla_{x}\langle f_{\varphi}(\mathbf{x}),\;g_{\psi}(\mathbf{y})\rangle=\nabla_{\mathbf{x}}\log\;p(\mathbf{y}|\mathbf{x})$. Therefore, we substitute this surrogate likelihood into Eq. 5) along with the DPS approximation which gives,
@@
$\mathcal{L}_{f\to t}=-\,\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p^{+}}\log\!\frac{\exp\!\big(\langle f_{\varphi}(\mathbf{x}),\,g_{\psi}(\mathbf{y})\rangle/\tau\big)}{\exp\!\big(\langle f_{\varphi}(\mathbf{x}),\,g_{\psi}(\mathbf{y})\rangle/\tau\big)+\mathbb{E}_{\mathbf{y}^{-}\sim p(\mathbf{y})}\!\left[\exp\!\big(\langle f_{\varphi}(\mathbf{x}),\,g_{\psi}(\mathbf{y}^{-})\rangle/\tau\big)\right]}.$ (9)
@@
$\mathcal{L}_{t\to f}=-\,\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p^{+}}\log\!\frac{\exp\!\big(\langle g_{\psi}(\mathbf{y}),\,f_{\varphi}(\mathbf{x})\rangle/\tau\big)}{\exp\!\big(\langle g_{\psi}(\mathbf{y}),\,f_{\varphi}(\mathbf{x})\rangle/\tau\big)+\mathbb{E}_{\mathbf{x}^{-}\sim p(\mathbf{x})}\!\left[\exp\!\big(\langle g_{\psi}(\mathbf{y}),\,f_{\varphi}(\mathbf{x}^{-})\rangle/\tau\big)\right]}.$ (10)
Adding Alignment Losses. To improve performance, we train the contrastive model with the symmetric losses in Eq. 9, 10 along with an additional alignment loss that further pull each matched floorplan–trajectory pair closer in the embedding space.
@@ 3.3 Improving Diffusion Inference
Intersection Penalty. We found it helpful to penalize intersections between walls and trajectories during inference. We add an intersection penalty $\mathcal{L}_{\text{intersect}}=\|\mathbf{y}\odot(1-\hat{\mathbf{x}}_{0})\|_{1}$ which counts (up to a constant scale) the total number of pixels where a trajectory overlaps a wall. Lowering this term during reverse diffusion steers updates toward wall–trajectory compatibility, yielding floorplans that respect the observed paths $\mathbf{y}$.
@@
Algorithm 1 CoGuide: Contrastive Likelihood Guidance for Spatial Inverse Problems
@@
6:   CoGuide: $G_{t}\leftarrow-\dfrac{1}{2\tau}\,\nabla_{\mathbf{x}_{t}}\!\left\|g_{\psi}(\mathbf{y})-f_{\varphi}(\hat{\mathbf{x}}_{0})\right\|_{2}^{2}+\lambda_{int}\|\mathbf{y}\odot(1-\hat{\mathbf{x}}_{0})\|_{1}$$\triangleright$ contrastive likelihood score

### v2

@@ Abstract
We consider a class of inverse problems characterized by forward operators that are partially specified, non-smooth, and non-differentiable. Although generative inverse solvers have made significant progress, we find that these forward operators introduce a distinct set of challenges. As a concrete instance, we consider the problem of reconstructing spatial layouts, such as floorplans, from human movement trajectories, where the underlying path-generation process is inherently non-differentiable and only partially known. In such problems, direct likelihood-based guidance becomes unstable, since the underlying path-planning process does not provide reliable gradients. We break-away from existing diffusion-based posterior samplers and reformulate likelihood-based guidance in a smoother embedding space. This embedding space is learned using a contrastive objective to bring compatible trajectory-floorplan pairs close together while pushing mismatched pairs apart. We show that this surrogate likelihood score in the embedding space provides a valid approximation to the true likelihood score, making it possible to steer the denoising process towards the posterior. Across extensive experiments, our model CoGuide produces more consistent reconstructions and is more robust than existing inverse-solvers and guided diffusion. Beyond spatial mapping, we show that our method can be applied more broadly, suggesting a route toward solving generalized blind inverse problems using diffusion models.
@@ 1 Introduction
Consider a user walking around in her home for a few minutes. Using some sensor, e.g., a smartphone, the user’s trajectory has been recorded. This trajectory is a sequence of location measurements inside the home, $\mathbf{y}=[y_{1},y_{2},\dots y_{n}]$, along which the user has walked. We ask, given this trajectory measurement, is it possible to infer the floorplan $\mathbf{x}$ of the home, where floorplan is the dimensions and layouts of the walls in the home. Observe that this is a spatial inverse problem, modeled as $\mathbf{y}=\mathcal{A}(\mathbf{x},\mathbf{n})$, because the way the human user walks, $\mathbf{y}$, is indeed a function of the layout of the home $\mathbf{x}$. This function is the $\mathcal{A}(.)$ operator, a policy in the user’s brain that plans the path from point A to point B, for a given floorplan $\mathbf{x}$. This path-planner is complex—it models factors such as distance walked, collision with walls and furniture, time to walk, number of turns, etc. Optimizations with such functions inherit this complexity since tiny changes in the floorplan—say a small hole in one wall—can drastically change the planned path. Hence, $\mathcal{A}(.)$ is in the regime of non-linear, non-differentiable, and partially observable operators, presenting a relatively new question (to the best of our knowledge) to diffusion-based inverse solvers. This paper concentrates on this specific problem of inverse floorplan estimation, but shows the potential to extend the approach to a broader family of spatial inverse problems.
@@
We train our Diffusion prior using public floorplan datasets; during inference, our method CoGuide generates floorplans for given (sparse, medium, or dense) trajectories. Results reliably outperform $6$ different baselines: $3$ that are augmentations of DPS with path-planners Yonetani et al. (2021); Kirilenko et al. (2023); Liu et al. (2024), $2$ that are established inverse solvers Zhu et al. (2023); Wang et al. (2024), and $1$ classifier free guidance (CFG) diffusion model Ho & Salimans (2022) that (over)fits to the joint distribution of trajectories and floorplans. We believe CoGuide has potential beyond this specific application of floorplan inference and demonstrate generalization through blind audio restoration (Appendix D), an inverse problem where the forward degradation operator is completely unknown. We discuss early thoughts on follow-on research directions to leverage contrastive learning in diffusion-based inverse solvers.
@@ 2 Preliminaries
Diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020); Song et al. (2020a; b) are a class of generative models capable of producing high-quality samples across a wide range of domains, including images Dhariwal & Nichol (2021), audio Kong et al. (2020b); Liu et al. (2023), video Ho et al. (2022), and 3D data Luo & Hu (2021); Poole et al. (2022). These models define generation as the reversal of a forward noising process that is formalized using a stochastic differential equation (SDE).
@@
Here, the prior score is easy to approximate by a trained diffusion model $s_{\theta}(\mathbf{x}_{t},t)$. The likelihood score, in contrast, is intractable as $\mathbf{y}$ depends only on $\mathbf{x}_{0}$, not directly on $\mathbf{x}_{t}$. In response to this, prior work such as Diffusion Posterior Sampling (DPS) Chung et al. (2022) have approximated the likelihood term as $p_{t}(\mathbf{y}|\mathbf{x}_{t})\approx p(\mathbf{y}|\hat{\mathbf{x}}_{0})$ where $\hat{\mathbf{x}}_{0}:=\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})=\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$ is obtained from a single denoising step using Tweedie’s formula: $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\mathbf{x}_{t}+\sigma_{t}^{2}\nabla_{\mathbf{x}_{t}}\log\;p_{t}(\mathbf{x}_{t})$
This formulation has been successfully used to guide diffusion models in solving a variety of inverse problems across scientific domains Zheng et al. (2025). CoGuide builds on this framework but makes a necessary departure owing to the challenges posed by the path-planning operator. We briefly discuss these operators before formulating the floorplan inference inverse problem.
Path Planners as Forward Operators. Unlike typical inverse problems (e.g., deblurring or inpainting), our measurements $\mathbf{y}$ are human-walked trajectories; thus, the forward operator $\mathcal{A}$ encodes a walking policy through an indoor layout (the floorplan $\mathbf{x}$). Directly modeling human navigation is difficult, so we use path-planners as proxies LaValle (2006); Hart et al. (1968); Gammell et al. (2015); Noreen et al. (2016). Empirical evidence indicates people favor short, direct routes with few turns Tong & Bode (2022) which aligns well with shortest-path planning. Therefore, the classical A* algorithm Hart et al. (1968) is a good choice for the forward operator. The floorplan is discretized into a grid graph, and given a start and end location, A* computes the shortest collision-free path, which we take as the predicted walking trajectory. To fit into the inverse problem framework, differentiable variants of A*, such as Neural A* Yonetani et al. (2021), TransPath Kirilenko et al. (2023), NRRT Wang et al. (2020), and Takahashi et al. (2019) are of interest; they all aim to enable gradient-based learning. Recent diffusion-based planners (DiPPeR Liu et al. (2024), PbDiff Luo et al. (2024)) further provide differentiable path generation. When inserted as forward operators, these planners often induce non-smooth objectives where small layout perturbations can cause large path changes, which in turn hinders convergence in gradient-based optimization. We analyze causes for this instability next and motivate our embedding-space likelihood surrogate.
@@ 3.1 Problem Formulation
Fig. 1 (Left) top row shows an unknown 2D floorplan $\mathbf{x}\in\mathbb{R}^{m\times n}$. A user walks in this floorplan (from point A to a destination point B, then from B to another destination C, and so on) and records a sequence of location measurements. The union of all these location measurements on a 2D image gives us the trajectories $\mathbf{y}\in\mathbb{R}^{m\times n}$ as shown in Fig. 1 (Left) bottom row. Given location sensors are noisy, the forward process is $\mathbf{y}=\mathcal{A}(\mathbf{x})+\mathbf{n}$ where $\mathcal{A}:\mathbb{R}^{m\times n}\!\to\!\mathbb{R}^{m\times n}$ approximates the human walk using a (generally nonlinear) A* path-planner, and $\mathbf{n}$ models additive Gaussian noise of the location sensor.
@@
Approximating the likelihood requires propagating $\hat{\mathbf{x}}_{0}$ through the $\mathcal{A}(.)$ operator and using its gradient to steer the diffusion prior $s_{\theta}(\mathbf{x}_{t},t)$; doing this stably lies at the heart of our problem.
@@ 3.2 Guidance through Contrastive Learning
where $\mathbf{x}\in\mathcal{X}$ and $\mathbf{y}\in\mathcal{Y}$ such that $\|f_{\varphi}(\mathbf{x})\|_{2}=\|g_{\psi}(\mathbf{y})\|_{2}=1$. We then use these encoders to define the likelihood-to-evidence ratio for a pair of inputs $(\mathbf{x},\mathbf{y})$ in the form of an un-normalized distribution $\pi$:
$\pi(\mathbf{y},\mathbf{x})\propto\exp(\langle f_{\varphi}(\mathbf{x}),\;g_{\psi}(\mathbf{y})\rangle\,/\,\tau)$ (7)
@@
Contrastive similarity as a likelihood surrogate. To see why this is a valid approach to approximating the true likelihood, we note that InfoNCE links contrastive learning to density estimation. Following Oord et al. (2018) we observe that the optimal contrastive classifier recovers a likelihood-to-evidence ratio. That is, when the InfoNCE loss attains its optimum, we get
$\exp(\langle f_{\varphi}(\mathbf{x}),g_{\psi}(\mathbf{y})\rangle\,/\,\tau)$ $\propto\frac{p(\mathbf{y}|\mathbf{x})}{p(\mathbf{y})}\implies\frac{1}{\tau}\langle f_{\varphi}(\mathbf{x}),\;g_{\psi}(\mathbf{y})\rangle=\log\;p(\mathbf{y}|\mathbf{x})-\log\;p(\mathbf{y})+C$ (8)
where $C$ is a constant independent of $\mathbf{x}$. Taking gradients with respect to $\mathbf{x}$ in Eq. 8 exactly recovers the likelihood score on the right-hand side i.e., $\frac{1}{\tau}\nabla_{x}\langle f_{\varphi}(\mathbf{x}),\;g_{\psi}(\mathbf{y})\rangle=\nabla_{\mathbf{x}}\log\;p(\mathbf{y}|\mathbf{x})$. Therefore, we substitute this surrogate likelihood into Eq. 5) along with the DPS approximation which gives,
@@
$\mathcal{L}_{f\to t}=-\,\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p^{+}}\log\!\frac{\exp\!\big(\langle f_{\varphi}(\mathbf{x}),\,g_{\psi}(\mathbf{y})\rangle/\tau\big)}{\exp\!\big(\langle f_{\varphi}(\mathbf{x}),\,g_{\psi}(\mathbf{y})\rangle/\tau\big)+\mathbb{E}_{\mathbf{y}^{-}\sim p(\mathbf{y})}\!\left[\exp\!\big(\langle f_{\varphi}(\mathbf{x}),\,g_{\psi}(\mathbf{y}^{-})\rangle/\tau\big)\right]}$ (10)
@@
$\mathcal{L}_{t\to f}=-\,\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p^{+}}\log\!\frac{\exp\!\big(\langle g_{\psi}(\mathbf{y}),\,f_{\varphi}(\mathbf{x})\rangle/\tau\big)}{\exp\!\big(\langle g_{\psi}(\mathbf{y}),\,f_{\varphi}(\mathbf{x})\rangle/\tau\big)+\mathbb{E}_{\mathbf{x}^{-}\sim p(\mathbf{x})}\!\left[\exp\!\big(\langle g_{\psi}(\mathbf{y}),\,f_{\varphi}(\mathbf{x}^{-})\rangle/\tau\big)\right]}$ (11)
Adding Alignment Losses. To improve performance, we train the contrastive model with the symmetric losses in Eq. 10, 11 along with an additional alignment loss that further pulls each matched floorplan–trajectory pair closer in the embedding space.
@@ 3.3 Improving Diffusion Inference
Intersection Penalty. We found it helpful to penalize intersections between walls and trajectories during inference. We add an intersection penalty $\mathcal{L}_{\text{intersect}}=\|\mathbf{y}\odot(1-\hat{\mathbf{x}}_{0})\|_{1}$, to all baselines and CoGuide, which counts (up to a constant scale) the total number of pixels where a trajectory overlaps a wall. Minimizing this penalty term during reverse diffusion steers updates toward wall–trajectory compatibility, yielding floorplans that respect the observed trajectories $\mathbf{y}$.
@@
Using conditional diffusion priors. (CFG+CoGuide) Using conditional diffusion priors, we further boost performance by combining a CFG-trained diffusion model $s_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$ with CoGuide so they can complement each other’s strengths. In this setup, CFG acts as a conditional prior over floorplans given trajectory measurements $\mathbf{c}$; the only change is that we replace CoGuide’s original diffusion prior with the CFG model. When the null input $\mathbf{c}=\varnothing$, we recover the original CoGuide behavior.
Algorithm 1 CoGuide: Contrastive Diffusion Guidance for Spatial Inverse Problems
@@
6:   CoGuide: $G_{t}\leftarrow-\dfrac{1}{2\tau}\,\nabla_{\mathbf{x}_{t}}\!\left\|g_{\psi}(\mathbf{y})-f_{\varphi}(\hat{\mathbf{x}}_{0})\right\|_{2}^{2}+\lambda_{int}\nabla_{\mathbf{x}_{t}}\|\mathbf{y}\odot(1-\hat{\mathbf{x}}_{0})\|_{1}$$\triangleright$ contrastive likelihood score